OpenStudy (anonymous):

A man on the fifth floor of a building shouts down to a person on the street. If the angle of elevation from the street to the man in the building is 35° and the man in the building is 40 feet up, about how far away from the building is the person on the street?

OpenStudy (anonymous):

sin35° =40/x x = 40/sin35 x= 40/.5736 x = 69.7350069735007